# Semi-classical imprints on quasinormal mode spectra

### v2

@@ IV.2 The wave equation
$r_{*}$ $=\int\!\dfrac{dr}{f(r)}\underset{r=r_{g}}{\approx}\frac{\log(r-r_{g})}{a_{1}}-\dfrac{a_{2}(r-r_{g})}{a_{1}^{2}}\ ,$ (33)
@@
$r_{*}=r+r_{g}\log\left(\frac{r-r_{g}}{r_{g}}\right)\ .$ (34)
@@
$\chi\equiv\-\frac{a_{2}(r-r_{g})}{a_{1}^{2}}\ .$ (36)
@@
$\phi(r)=e^{-i\omega\chi}(r-r_{g})^{-\tfrac{i\omega}{a_{1}}}e^{2i\omega r}\left(\frac{r}{r_{g}}\right)^{\!ir_{g}\omega}r^{\tfrac{i\omega}{a_{1}}}R(r)$ (38)
@@ V Matrix method
Thus we modify the algorithm to proceed in a sequence of steps, as follows. Since our interpolating metric (31) and radial wavefunction (39) possess smooth limits to the Schwarzschild case when $a_{1}\rightarrow 1/r_{g}$ and $a_{2}\rightarrow-1/r_{g}^{2}$, we proceed in a sequence of steps beginning with this limit, then increment the coefficients $a_{1}$ and $a_{2}$ in finite steps $\Delta a_{1}$ and $\Delta a_{2}$ until the target values are reached. At each step, the root finder is seeded with the result of the previous step and picks out the solution which minimizes the $L2$ distance from the quasinormal frequency computed in the previous step. In this way the real quasinormal frequency is identified and tracked, which should vary smoothly away from its Schwarzschild value as the interpolating metric does.
@@ VI Quasinormal modes
We now compute the low-lying $l=0,1,2$ quasinormal frequencies for the fundamental modes. Note that in spherical symmetry modes of the same $l$ but different $m$ are degenerate. In the absence of observational constraints on scalar perturbations, we assume a hypothetical 10% bound on the deviations of these frequencies from their Schwarzschild values, and use these to infer bounds and sensitivity on the leading coefficients $(a_{1},a_{2})$ of (14) and thus, the leading coefficients $(E,e_{2})$ of the EMT expansion near the horizon. For all computations, a sufficiently high number of grid points is chosen to achieve a numerical error of at most $\mathcal{O}(10^{-4})$, below which we expect the choice of interpolation procedure and truncation of the series (29) to dominate the error.
@@ VI.1 $l=0$
We compute the frequency and damping time of the (000) mode on a uniformly spaced grid with $N=30$, as a function of $a_{1}$ with $a_{2}$ fixed and $a_{2}$ with $a_{1}$ fixed. The results are displayed in Figure4 and Figure5. For the Schwarzschild metric, the scalar mode has frequency $\omega_{000}=0.220952-0.209795i$, with a 10% deviation implying
@@
Figure 4: Deviation from GR prediction for the frequency and damping time of (000) as a function of $a_{1}$ with $a_{2}=0$. The red shaded region corresponds to a 10% deviation from the Schwarzschild value as given in (VI.1). The step size is $\Delta a_{1}=0.000025$ in all panels.
Figure 5: Deviation from GR prediction for the frequency and damping time of (000) as a function of $a_{2}$ with $a_{1}=0$. The red shaded region corresponds to a 10% deviation from the Schwarzschild value as given in (VI.1). The step size is $\Delta a_{2}=0.000001$ in all panels.
In both the frequency and damping time of the $(000)$ mode, fractional deviations from the Schwarzschild frequency on the order $0.1$ occur already with $\Delta a_{1}=10^{-4}$, showing high sensitivity to the leading coefficient of (14). As a check, we also compare the result obtained for $\omega_{000}$ using the interpolating metric (31) with the value obtained using (32) on the same grid. For example, taking $a_{1}=0.9999$ with $a_{2}=-1$ we find
$\omega_{000}$ $=0.22080-0.21006i\qquad\text{with}\quad f(r)$ (52)
$\omega_{000}$ $=0.22167-0.21567i\qquad\text{with}\quad g(r)\ .$ (53)
@@ VI.2 $l=1$
Next we compute the $l=1$ fundamental mode for various values of $a_{1}$ and $a_{2}$, and again consider a hypothetical 10% deviation from the Schwarzschild quasinormal frequency $\omega_{100}=0.967287-0.193518i$. The results are shown in Figure6 and Figure7.
Figure 6: Deviation from GR prediction for the frequency and damping time of (100) as a function of $a_{1}$ with $a_{2}=0$. The red shaded region corresponds to a 10% deviation from the Schwarzschild value as given in (VI.1). The step size is $\Delta a_{1}=0.00002$ in all panels.
Figure 7: Deviation from GR prediction for the frequency and damping time of (100) as a function of $a_{2}$ with $a_{1}=0$. The red shaded region corresponds to a 10% deviation from the Schwarzschild value as given in (VI.1). The step size is $\Delta a_{2}=0.000002$ in all panels.
@@
The above observation suggests that with finite precision, a single measurement of the $(100)$ mode is insufficient to distinguish between a rotating Kerr black hole and a spherically symmetric semi-classical black hole in cases where $a_{1}<1$. When $a_{1}>1$ (and $a_{2}<-1$) the quasinormal frequency and damping time instead increase relative to the Schwarzschild limit and the mode instead mimics the $l=1,m=1$ Kerr counterpart. Notably, a 10% deviation in $\omega_{100}$ is achieved by a deviation of just $|\Delta a_{1}|=0.0001$ relative to the Schwarzschild value, while a comparable shift in $\omega_{100}$ due to spin requires $|\Delta a|\approx 0.5$.
@@ VI.3 $l=2$
Figure 8: Deviation from GR prediction for the frequency and damping time of (200) as a function of $a_{1}$ with fixed $a_{2}=0$. The red shaded region corresponds to a 10% deviation from the Schwarzschild value. The step size is $\Delta a_{1}=0.00002$ in all panels.
Figure 9: Deviation from GR prediction for the frequency and damping time of (200) as a function of $a_{2}$ with fixed $a_{1}=0$. The red shaded region corresponds to a 10% deviation from the Schwarzschild value. The step size is $\Delta a_{2}=0.000002$ in all panels.
All modes display similar sensitivity to underlying changes in $(a_{1},a_{2})$, with the highest sensitivity to $a_{1}$ occurring for the frequency of the $(000)$ mode, and the highest sensitivity for $a_{2}$ being tied to the damping time of $(000)$. A hypothetical maximum deviation of 10% in their respective quasinormal frequencies translates to a constraint on the leading components of the EMT near the apparent horizon as
@@ VII Discussion
We find that deviations in the leading coefficient $a_{1}$ of the order $10^{-4}$ and of the subleading coefficient $a_{2}$ of order $10^{-5}$ lead to a deviation in the $l\in\{0,1,2\}$ fundamental modes exceeding 10% of their Schwarzschild values, implying that tight constraints on deviations from the Schwarzschild solution may arise from semi-classical effects near the horizon. Notably, we show that spherically symmetric variations of the near-horizon geometry away from the Schwarzschild metric are able to mimic the imprint that non-zero angular momentum would have on certain scalar quasinormal modes. In principle such an effect has the potential to obscure the distinction between a Kerr black hole and a non-rotating black hole with quantum effects near the horizon in the absence of additional independent measurements, highlighting the importance of multi-modal measurements in distinguishing between various black hole mimickers.
The advantage of our scheme is its relative model-independence. However, one limitation of the two-point Padé interpolation lies in the existence of poles in the interpolating function for some (potentially interesting) values of the metric parameters. For example, both functions $f(r)$ and $g(r)$ cannot accommodate the values of $(a_{1},a_{2})$ that correspond to the Bardeen metric. This is a drawback that our M-approximants share with the C-approximants used in Ref. [22]. We have shown that in the example we study here, the variation in QNM frequencies from different valid (pole-free) interpolations is appropriately small. We have also considered scalar perturbations in spherical symmetry only, and only in the frequency domain. Finally, while the semi-classical near-horizon metric (14) describes a wide variety of black hole models in spherical symmetry—such as Reissner-Nordström and Bardeen regular black holes–the interpolation (IV) cannot support these geometries without additional assumptions on the terms entering the asymptotic expansion of the metric.
Future work will will proceed in several directions. First by generalizing to gravitational perturbations, which will be necessary to make concrete comparisons to observational data. We will also aim to develop a more sophisticated interpolation scheme which can incorporate a wider class of metrics, including boundary conditions consistent with the presence of a cosmological constant. Finally we aim to generalize our work to dynamical spacetimes, and compute QNMs for slowly evolving semi-classical black holes in a self-consistent setting.
@@ Appendix A The semi-classical metric
where $c_{1}=8\pi r_{g}^{2}E$ (and thus $f=(1-8\pi Er_{g}^{2})/r_{g}\tilde{x}+\mathcal{O}(\tilde{x}^{2})$), the coefficient $h_{1}=8(e_{2}+p_{2})r_{0}^{2}/(1-c_{1})^{2}$ is determined by the coefficients of the second order terms in the EMT components, etc. If the black hole has no hair whatsoever, then $h\equiv 0$, while for short hair it is a rapidly vanishing function of $\tilde{x}$. The next coefficient in the expansion of $f$ is
$a_{2}$ $=\frac{1}{3\,r_{g}^{2}\left(8\pi E\,r_{g}^{2}-1\right)^{3}}\bigg{[}192\pi^{2}E\,^{2}r_{g}^{4}\left(4\pi e_{2}r_{g}^{4}+3\right)$
@@
Appendix B Continued fraction representation
The M-fraction Padé expansion of a function with simultaneous expansions (29) and (20) is given by adapting Eq. (24). For $n=3$ it is
$M_{3}(\tilde{x})=1-\cfrac{1}{1+\dfrac{\tilde{x}}{r_{g}}+\cfrac{\left(a_{1}-\dfrac{1}{r_{g}}\right)\tilde{x}}{1-\left(a_{1}-\dfrac{1}{r_{g}}\right)\tilde{x}+\cfrac{\left(\dfrac{1-a_{2}r_{g}^{2}-2a_{1}r_{g}}{r_{g}\left(a_{1}r_{g}-1\right)}\right)\tilde{x}}{1-\left(\dfrac{1-a_{2}r_{g}^{2}-2a_{1}r_{g}}{r_{g}\left(a_{1}r_{g}-1\right)}\right)\tilde{x}}}}$ (72)
@@ Appendix B Continued fraction representation
$\bar{A}$ $=1-a_{2}r_{g}^{2}-2a_{1}r_{g},\qquad\bar{B}=a_{2}r_{g}^{3}+a_{1}r_{g}^{2},$
$\bar{C}$ $=a_{1}r_{g}^{3},\qquad\bar{D}=0,$
$\bar{\bar{B}}$ $=r_{g}-a_{1}r_{g}^{2},\qquad\bar{\bar{C}}=r_{g}^{2},\qquad\bar{\bar{D}}=r_{g}^{3}.$
@@
$a_{1}r_{g}=1+{a^{C}_{1}}-2\epsilon,\qquad a_{2}r_{g}^{2}=-1-4a^{C}_{1}-a^{C}_{1}a^{C}_{2}+5\epsilon\ .$ (76)
@@
Requiring that these poles do not occur outside of the horizon determines the domain of validity of the approximation in the parameter space of the metric. E.g., for the Bardeen metrics, direct evaluation gives that
$a_{2}^{C}=-\frac{96+83\lambda^{2}}{64+42\lambda^{2}}<1\ ,$ (79)

### v3

@@ IV.2 The wave equation
$r_{*}$ $=\int\!\dfrac{dr}{f(r)}{\approx}\frac{r_{g}}{\alpha_{1}}\log(r-r_{g})-\dfrac{\alpha_{2}(r-r_{g})}{\alpha_{1}^{2}}\ ,$ (33)
@@
$r_{*}=r+r_{g}\log\left(\frac{\tilde{x}}{r_{g}}\right)\ .$ (34)
@@
$\chi{\vcentcolon=}\-\frac{\alpha_{2}(r-r_{g})}{\alpha_{1}^{2}}\ .$ (36)
@@
$\phi(r)=e^{-i\omega\chi}(r-r_{g})^{-\tfrac{i\omega}{a_{1}}}e^{2i\omega r}\left(\frac{r}{r_{g}}\right)r^{\tfrac{i\omega}{a_{1}}}R(r)$ (38)
@@ V Matrix method
Thus we modify the algorithm to proceed in a sequence of steps, as follows. Since our interpolating metric (31) and radial wavefunction (39) possess smooth limits to the Schwarzschild case when $\alpha_{1}\rightarrow 1$ and $\alpha_{2}\rightarrow-1$, we proceed in a sequence of steps beginning with this limit, then increment the coefficients $\alpha_{1}$ and $\alpha_{2}$ in finite steps $\Delta\alpha_{1}$ and $\Delta\alpha_{2}$ until the target values are reached. At each step, the root finder is seeded with the result of the previous step and picks out the solution which minimizes the $L2$ distance from the quasinormal frequency computed in the previous step. In this way the real quasinormal frequency is identified and tracked, which should vary smoothly away from its Schwarzschild value as the interpolating metric does.
@@ VI Quasinormal modes
We now compute the low-lying $l=0,1,2$ quasinormal frequencies for the fundamental modes. Note that in spherical symmetry modes of the same $l$ but different $m$ are degenerate. In the absence of observational constraints on scalar perturbations, we assume a hypothetical 10% bound on the deviations of these frequencies from their Schwarzschild values, and use these to infer bounds and sensitivity on the leading coefficients $(\alpha_{1},\alpha_{2})$ of (14) and thus, the leading coefficients $(E,e_{2})$ of the EMT expansion near the horizon. For all computations, a sufficiently high number of grid points is chosen to achieve a numerical error of at most $\mathcal{O}(10^{-4})$, below which we expect the choice of interpolation procedure and truncation of the series (29) to dominate the error.
@@ VI.1 $l=0$
We compute the frequency and damping time of the (000) mode on a uniformly spaced grid with $N=30$, as a function of $\alpha_{1}$ with $\alpha_{2}$ fixed and $\alpha_{2}$ with $\alpha_{1}$ fixed. The results are displayed in Figure5 and Figure6. For the Schwarzschild metric, the scalar mode has frequency $\omega_{000}=0.220952-0.209795i$, with a 10% deviation implying
@@
Figure 5: Deviation from GR prediction for the frequency and damping time of (000) as a function of $\alpha_{1}$ with $\alpha_{2}=0$. The red shaded region corresponds to a 10% deviation from the Schwarzschild value as given in (VI.1). The step size is $\Delta\alpha_{1}=0.000025$ in all panels.
Figure 6: Deviation from GR prediction for the frequency and damping time of (000) as a function of $\alpha_{2}$ with $\alpha_{1}=0$. The red shaded region corresponds to a 10% deviation from the Schwarzschild value as given in (VI.1). The step size is $\Delta\alpha_{2}=0.000001$ in all panels.
In both the frequency and damping time of the $(000)$ mode, fractional deviations from the Schwarzschild frequency on the order $0.1$ occur already with $\Delta\alpha_{1}=10^{-4}$, showing high sensitivity to the leading coefficient of (14). As a check, we also compare the result obtained for $\omega_{000}$ using the interpolating metric (31) with the value obtained using (32) on the same grid. For example, taking $\alpha_{1}=0.9999$ with $\alpha_{2}=-1$ we find
$\omega_{000}$ $=0.22080-0.21006i\qquad\text{with}\quad f_{5}(r)$ (52)
$\omega_{000}$ $=0.22167-0.21567i\qquad\text{with}\quad f_{4}(r)\ .$ (53)
@@ VI.2 $l=1$
Next we compute the $l=1$ fundamental mode for various values of $\alpha_{1}$ and $\alpha_{2}$, and again consider a hypothetical 10% deviation from the Schwarzschild quasinormal frequency $\omega_{100}=0.967287-0.193518i$. The results are shown in Figure7 and Figure8.
Figure 7: Deviation from GR prediction for the frequency and damping time of (100) as a function of $\alpha_{1}$ with $\alpha_{2}=0$. The red shaded region corresponds to a 10% deviation from the Schwarzschild value as given in (VI.1). The step size is $\Delta a_{1}=0.00002$ in all panels.
Figure 8: Deviation from GR prediction for the frequency and damping time of (100) as a function of $\alpha_{2}$ with $\alpha_{1}=0$. The red shaded region corresponds to a 10% deviation from the Schwarzschild value as given in (VI.1). The step size is $\Delta\alpha_{2}=0.000002$ in all panels.
@@
The above observation suggests that with finite precision, a single measurement of the $(100)$ mode is insufficient to distinguish between a rotating Kerr black hole and a spherically symmetric semi-classical black hole in cases where $a_{1}<1$. When $\alpha_{1}>1$ (and $\alpha_{2}<-1$) the quasinormal frequency and damping time instead increase relative to the Schwarzschild limit and the mode instead mimics the $l=1,m=1$ Kerr counterpart. Notably, a 10% deviation in $\omega_{100}$ is achieved by a deviation of just $|\Delta\alpha_{1}|=0.0001$ relative to the Schwarzschild value, while a comparable shift in $\omega_{100}$ due to spin requires $|\Delta\alpha_{1}|\approx 0.5$.
@@ VI.3 $l=2$
Figure 9: Deviation from GR prediction for the frequency and damping time of (200) as a function of $\alpha_{1}$ with fixed $a_{2}=0$. The red shaded region corresponds to a 10% deviation from the Schwarzschild value. The step size is $\Delta\alpha_{1}=0.00002$ in all panels.
Figure 10: Deviation from GR prediction for the frequency and damping time of (200) as a function of $\alpha_{2}$ with fixed $\alpha_{1}=0$. The red shaded region corresponds to a 10% deviation from the Schwarzschild value. The step size is $\Delta\alpha_{2}=0.000002$ in all panels.
All modes display similar sensitivity to underlying changes in $(\alpha_{1},\alpha_{2})$, with the highest sensitivity to $\alpha_{1}$ occurring for the frequency of the $(000)$ mode, and the highest sensitivity for $\alpha_{2}$ being tied to the damping time of $(000)$. A hypothetical maximum deviation of 10% in their respective quasinormal frequencies translates to a constraint on the leading components of the EMT near the apparent horizon as
@@ VII Discussion
We find that deviations in the leading coefficient $\alpha_{1}$ of the order $10^{-4}$ and of the subleading coefficient $\alpha_{2}$ of order $10^{-5}$ lead to a deviation in the $l\in\{0,1,2\}$ fundamental modes exceeding 10% of their Schwarzschild values, implying that tight constraints on deviations from the Schwarzschild solution may arise from semi-classical effects near the horizon. Notably, we show that spherically symmetric variations of the near-horizon geometry away from the Schwarzschild metric are able to mimic the imprint that non-zero angular momentum would have on certain scalar quasinormal modes. In principle such an effect has the potential to obscure the distinction between a Kerr black hole and a non-rotating black hole with quantum effects near the horizon in the absence of additional independent measurements, highlighting the importance of multi-modal measurements in distinguishing between various black hole mimickers.
The advantage of our scheme is its relative model-independence. However, one limitation of the two-point Padé interpolation lies in the existence of poles in the interpolating function for some (potentially interesting) values of the metric parameters. For example, both functions $f_{5}(r)$ and $f_{4}(r)$ cannot accommodate the values of $(\alpha_{1},\alpha_{2})$ that correspond to the Bardeen metric. This is a drawback that our M-approximants share with the C-approximants used in Ref. [22]. We have shown that in the example we study here, the variation in QNM frequencies from different valid (pole-free) interpolations is appropriately small. We have also considered scalar perturbations in spherical symmetry only, and only in the frequency domain. Finally, while the semi-classical near-horizon metric (14) describes a wide variety of black hole models in spherical symmetry—such as Reissner-Nordström and Bardeen regular black holes–the interpolation (IV) cannot support these geometries without additional assumptions on the terms entering the asymptotic expansion of the metric.
Future work will proceed in several directions. First by generalizing to gravitational perturbations, which will be necessary to make concrete comparisons to observational data. We will also aim to develop a more sophisticated interpolation scheme which can incorporate a wider class of metrics, including boundary conditions consistent with the presence of a cosmological constant. Finally we aim to generalize our work to dynamical spacetimes, and compute QNMs for slowly evolving semi-classical black holes in a self-consistent setting. In each case it is desirable to explore as much of the parameter space describing the near-horizon metric as possible, and important classes of problems can only be studied when an interpolating metric can be constructed. Likely a combination of the M-approximants and the C-approximants offers the best solution, and future work aims to develop even more flexible interpolation schemes.
@@ Appendix A The semi-classical metric
where $c_{1}=8\pi r_{g}^{2}E$ (and thus $f=(1-8\pi Er_{g}^{2})\tilde{x}/r_{g}+\mathcal{O}(\tilde{x}^{2})$), the coefficient $h_{1}=8(e_{2}+p_{2})r_{0}^{2}/(1-c_{1})^{2}$ is determined by the coefficients of the second order terms in the EMT components, etc. If the black hole has no hair whatsoever, then $h\equiv 0$, while for short hair it is a rapidly vanishing function of $\tilde{x}$. The next coefficient in the expansion of $f$ is
$\alpha_{2}$ $=\frac{1}{3\left(8\pi E\,r_{g}^{2}-1\right)^{3}}\bigg{[}192\pi^{2}E\,^{2}r_{g}^{4}\left(4\pi e_{2}r_{g}^{4}+3\right)$
@@
Appendix B Continued fraction representations
This Appendix compares the M-fraction approximation used in this work to the continued fraction parameterization of Rezzolla and Zhidenko (RZ) [22]. The two schemes both involve continued fractions but are distinct in a number of important ways. The RZ construction involves parameterizing, or approximating, a known metric by truncating a continued fraction $\tilde{A}(x)$ (Eq. (19a) in [22]) to a given order. In so doing, one obtains an approximate metric with some pre-determined near-horizon and asymptotic form which can be revealed by expanding the metric near $r=r_{g}$ and $r=\infty$. The order of approximation in $\tilde{A}(x)$ fixes both of these, and the coefficients $a_{i}$ defining the continued fraction $\tilde{A}(x)$ appear in both expansions at some finite order.
By contrast, our scheme involves choosing independently the near-horizon and asymptotic forms of an unknown metric first, by specifying some finite number of terms in its near and far expansion. The M-fraction then allows one to construct a metric which interpolates between the two expansions, and matches them simultaneously up to the specified order in both the near and far regions. Our construction allows one to specify the form of these expansions independently, whereas in the RZ scheme, the coefficients $\{a_{i}\}$ of the continued fraction appear in the coefficients of both the near and asymptotic expansions of the resulting metric. Modifying one of the terms in the near-expansion (for example by changing $a_{2}$) necessarily also changes the metric in the far region, as $a_{2}$ appears in those coefficients as well. Thus, the RZ construction does not allow for the independent variation of the near-horizon and asymptotic regions, while the M-fraction construction presented here does. This is desirable for studying deviations of the near-horizon geometry of known black hole solutions without necessarily requiring that such deviations propagate (or can be observed) in the asymptotic region.
Figure 11: Comparison of the lowest-order M-fraction (yellow) and Rezzolla–Zhidenko (red) approximations of the Bardeen metric (72) with $q=0.01$. Top: The shaded region corresponds to values of the expansion coefficients ($\alpha_{1},\alpha_{2}$) for which the given approximating metric diverges outside of the horizon. Bottom: A plot of the approximating metric function for the metric given by the black point.
In Figure 11, we depict the region of validity of the lowest-order non-trivial M-fraction and RZ approximations for the Bardeen metric:
$f(r)=1-\frac{r^{2}\left(r_{g}^{2}+q^{2}\right)^{3/2}}{r_{g}^{2}\left(r^{2}+q^{2}\right)^{3/2}}$ (72)
Note the complementary nature of the two approximations. Each point in the figure represents a metric which describes a small deviation away from the Bardeen form, as parameterized by deviations in the leading near-horizon and asymptotic expansion coefficients. For each of the two approximations, the shaded region corresponds to a choice of coefficients which leads to a divergence in the metric function outside of the horizon, rendering the approximation unusable. This is demonstrated by the metric described by the black point in the top panel.
As a simple example, we illustrate the relationship between the parameters of the two schemes. The M-fraction Padé expansion of a function with simultaneous expansions (29) and (20) is given by adapting Eq. (24). For $n=3$ it is
$M_{3}(\tilde{x})=1-\cfrac{1}{1+\dfrac{1}{r_{g}}\,\tilde{x}+\cfrac{\left(\dfrac{\alpha_{1}-{1}}{r_{g}}\right)\tilde{x}}{1-\left(\dfrac{\alpha_{1}-1}{r_{g}}\right)\tilde{x}+\cfrac{\left(\dfrac{1-\alpha_{2}-2\alpha_{1}}{r_{g}\left(\alpha_{1}-1\right)}\right)\tilde{x}}{1-\left(\dfrac{1-\alpha_{2}-2\alpha_{1}}{r_{g}\left(\alpha_{1}-1\right)}\right)\tilde{x}}}}$ (73)
@@ Appendix B Continued fraction representations
$\bar{A}$ $=(1-2\alpha_{1}-\alpha_{2})$
$\bar{B}$ $=(\alpha_{1}r_{g}+\alpha_{2}r_{g})$
$\bar{C}$ $=\alpha_{1}r_{g}^{2}$
$\bar{D}$ $=0$
$\bar{\bar{B}}$ $=(r_{g}-\alpha_{1}r_{g})$
$\bar{\bar{C}}$ $=r_{g}^{2}$
$\bar{\bar{D}}$ $=r_{g}^{3}\ .$
@@
$\alpha_{1}=1+{a^{C}_{1}}-2\epsilon,\qquad\alpha_{2}=-1-4a^{C}_{1}-a^{C}_{1}a^{C}_{2}+5\epsilon\ .$ (77)
@@
Requiring that these poles do not occur outside of the horizon determines the domain of validity of the approximation in the parameter space of the metric. For the Bardeen metrics, direct evaluation gives that
$a_{1}^{C}=-\tfrac{3}{2}\lambda^{2}+\mathcal{O}\big{(}\lambda^{4}\big{)}\ ,$ (80)
and
$a_{2}^{C}=-\frac{3}{2}-\frac{5}{16}\lambda^{2}+\mathcal{O}\big{(}\lambda^{4}\big{)}<1\ ,$ (81)